The amount of time it takes an asteroid, whose average distance from the Sun is 4.70 astronomical units, to complete one orbit around the Sun is ____ years
Answer #1

For Earth the average distance(R) from sun is 1 AU and the period(T) is 1 year

From Kepler's third law

T2 R3 T R3/2

Hence T2 / R3 (for asteroid)= constant = T2 / R3 (for earth) = 1

So T = R3/2

T = 4.73/2 = 10.19 years